# On polynomial automorphisms of Nagata type

### v1

@@ 3. Statement of the main results
(9) $\varphi=\varphi^{2n}+\varphi^{2n-1}z+\cdots+\varphi^{1}z^{2n-1}+\varphi^{0}z^{2n}$
with $\varphi^{2n-k}$ be a homogeneous polynomial in $F[x,y]$ of degree $2n-k$ for $0\leq k\leq 2n$. By Lemma 4, we get the following equalities:
(10) $\varphi^{2n}$ $=ay^{2n}\,\,\,\,\textnormal{and}\,\,\,\,\varphi^{1}=bx$
(11) $2y\varphi^{2n-k-1}_{x}$ $=\varphi^{2n-k}_{y}\,\,\,\textnormal{para todo}\,\,\,k=0,1,\dots,2n-2\,.$
We claim that there are constants $a,a^{2n-k}_{j}$, with $1\leq k\leq n$ y $10\leq j\leq k-$ 1, such that:
(12) $\varphi^{2n-k}$ $=a\binom{n}{k}y^{2n-2k}x^{k}+\sum_{j=0}^{k-1}a^{2n-k}_{k-1-j}y^{2n-2k+1+j}x^{k-1-j}$
Indeed, taking $k=0$, from these equalities (10) and (11) we find that:
$2y\varphi^{2n-1}_{x}=\varphi^{2n}_{y}=2nay^{2n-1}$
so $\varphi_{x}^{2n-1}=nay^{2n-2}$. Hence, by integrating with respect to $x$, we obtain
$\varphi^{2n-1}=nay^{2n-2}x+a^{2n-1}_{0}y^{2n-1}\,,$
@@
$\varphi^{2n-1}=a\binom{n}{1}y^{2n-2}x+a^{2n-1}_{0}y^{2n-1}\,,$
thus (12) is true for $k=1$. Suppose we have (12). Taking the derivative with respect to $y$, we have
$\varphi^{2n-k}_{y}=2a\binom{n}{k}(n-k)y^{2n-2k-1}x^{k}+\sum_{j=0}^{k-1}\,.(2n-2k+1+j)a^{2n-k}_{k-1-j}y^{2n-2k+j}x^{k-1-j}$
@@
$\varphi^{2n-k-1}_{x}=a\binom{n}{k}(n-k)y^{2n-2k-2}x^{k}+\sum_{j=0}^{k-1}\frac{(2n-2k+1+j)a^{2n-k}_{k-1-j}}{2}y^{2n-2k-1+j}x^{k-1-j}\,.$
Integrating with respect to $x$, there is a constant $a^{2n-k-1}_{0}\in F$ such that
$\varphi^{2n-k-1}=a\binom{n}{k+1}y^{2n-2k-2}x^{k+1}+\sum_{j=0}^{k-1}\frac{(2n-2k+1+j)a^{2n-k}_{k-1-j}}{2(k-1-j)}y^{2n-2k-1+j}x^{k-j}+a^{2n-k-1}_{0}y^{2n-k-1}\,.$
Defining $a^{2n-k-1}_{k-j}=\frac{(2n-2k+1+j)a^{2n-k}_{k-1-j}}{2(k-1-j)}$, we get
$\varphi^{2n-k-1}=a\binom{n}{k+1}y^{2n-2k-2}x^{k+1}+\sum_{j=0}^{k}a^{2n-k-1}_{k-j}y^{2n-2k-1+j}x^{k-j}$
as wished, and this completes the recurrence. Now, replacing the expressions obtained in (10) and (12) in the equation (9), we have
$\varphi$ $=\sum_{k=0}^{n}\varphi^{2n-k}z^{k}+\sum_{k=n+1}^{2n}\varphi^{2n-k}z^{k}$
$=\varphi^{2n}+\sum_{k=1}^{n}\varphi^{2n-k}z^{k}+\sum_{k=n+1}^{2n}\varphi^{2n-k}z^{k}$
$=ay^{2n}+\sum_{k=1}^{n}\left(a\binom{n}{k}y^{2n-2k}x^{k}+\sum_{j=0}^{k-1}a^{2n-k}_{k-1-j}y^{2n-2k+1+j}x^{k-1-j}\right)z^{k}+\sum_{k=n+1}^{2n}\varphi^{2n-k}z^{k}$
$=a\sum_{k=0}^{n}\binom{n}{k}y^{2n-2k}x^{k}z^{k}+\sum_{k=1}^{n}\left(\sum_{j=0}^{k-1}a^{2n-k}_{k-1-j}y^{2n-2k+1+j}x^{k-1-j}\right)z^{k}+\sum_{k=n+1}^{2n}\varphi^{2n-k}z^{k}\,.$
From the last equality, we deduce that
$\varphi=a(xz+y^{2})^{n}+z\psi,$
where $\psi\in F[x,y,z]$ is a homogeneous polynomial of degree equal to $12n-$ 1. By the linearity of the solutions of the equation (5), we obtain that $z\psi$ is a solution, hence $\psi$ is also a solution. ($b$) Let us suppose that:
(13) $\varphi=\varphi^{2n+1}+\varphi^{2n}+\cdots+\varphi^{1}z^{2n}+\varphi^{0}z^{2n+1}\,,$
where $\varphi^{2n-k}$ is a homogeneous polynomial in $F[x,y]$ of degree equal to $2n+1-k$, for all $10\leq k\leq 2n+$ 1. By Lemma 4, we get the following equalities:
(14) $\varphi^{2n+1}$ $=ay^{2n+1}$
(15) $2y\varphi^{2n-k}_{x}$ $=\varphi^{2n+1-k}_{y}\,\,\,\textnormal{para todo}\,\,\,k=0,1,\dots,2n-1\,.$
As in the even case, we claim that there are constants $a,a_{j}^{2n-k}$, with $10\leq k\leq n-$ 1 and $0\leq j\leq k$, such that:
(16) $\varphi^{2n-k}=\frac{(2n-2k+1)\cdots(2n-1)(2n+1)}{2^{k}\cdot(k+1)!}ay^{2n-2k-1}x^{k+1}+\sum_{j=0}^{k}b^{2n-k}_{k-j}y^{2n-2k+j}x^{k-j}\,.$
Replacing $k=0$ in the equation (14) and joining with (15), we have
$2y\varphi^{2n}_{x}=\varphi^{2n+1}_{y}=(2n+1)ay^{2n}\,,$
hence $\varphi_{x}^{2n}=(2n+1)ay^{2n-1}$. Integrating with respect to $x$, we obtain
$\varphi^{2n}=(2n+1)ay^{2n-1}x+b^{2n}_{0}y^{2n}$
thus (16) is true for $k=1$. Suppose we have (16). Taking the derivative with respect to $y$, we have
$\varphi^{2n-k}_{y}=\frac{(2n-2k-1)\cdots(2n-1)(2n+1)}{2^{k}\cdot(k+1)!}ay^{2n-2k-2}x^{k+1}+\sum_{j=0}^{k}(2n-2k+j)b^{2n-k}_{k-j}y^{2n-2k+j-1}x^{k-j}\,.$
From the equality (15), we get
$\varphi^{2n-k-1}_{x}=\frac{(2n-2k-1)\cdots(2n-1)(2n+1)}{2^{k+1}\cdot(k+1)!}ay^{2n-2k-3}x^{k+1}+\sum_{j=0}^{k}\frac{(2n-2k+j)b^{2n-k}_{k-j}}{2}y^{2n-2k+j-2}x^{k-j}\,.$
Integrating with respect to $x$, there is a constant $b^{2n-k-1}_{0}\in F$ such that $\varphi^{2n-k-1}$ is equal to
$\frac{(2n-2k-1)\cdots(2n-1)(2n+1)}{2^{k+1}\cdot(k+2)!}ay^{2n-2k-3}x^{k+2}+\sum_{j=0}^{k}\frac{(2n-2k+j)b^{2n-k}_{k-j}}{2(k-j+1)}y^{2n-2k+j-2}x^{k-j+1}+b^{2n-k-1}_{0}y^{2n-k-1}\,.$
Defining $b^{2n-k-1}_{k-j}=\frac{(2n-2k+j)b^{2n-k}_{k-j}}{2(k-j+1)}$, we get
$\varphi^{2n-k-1}=\frac{(2n-2k+-1)\cdots(2n-1)(2n+1)}{2^{k+1}\cdot(k+2)!}ay^{2n-2k-3}x^{k+2}+\sum_{j=0}^{k+1}\frac{(2n-2k+j)b^{2n-k}_{k-j}}{2(k-j+1)}y^{2n-2k+j-2}x^{k-j+1}$
as wished, and this completes the recurrence. In particular, replacing $k=n-1$ in (16), we obtain:
$\varphi^{n+1}=\frac{1\cdot 3\cdots(2n+1)}{2^{n-1}\cdot n!}ayx^{n}+\sum_{j=0}^{n-1}b^{n+1}_{n-1-j}y^{j+2}x^{n-1-j}\,.$
We solve the equation (15) for $k=n$ and we get:
$-2y\varphi^{n}_{x}$ $=\varphi^{n+1}_{y}\,,$
(17) $-2y\varphi^{n}_{x}$ $=\frac{1\cdot 3\cdots(2n+1)}{2^{n-1}\cdot n!}ax^{n}+\sum_{j=0}^{n-1}(j+2)b^{n+1}_{n-1-j}y^{j+1}x^{n-1-j}\,.$
From the equality (17), we deduce that $a=0$ and as consequence from this we obtain that $\varphi^{2n+1}=0$. Therefore,
$\varphi=z\varphi^{2n}+z^{2}\varphi^{2n-1}+\cdots+z^{2n}\varphi^{1}+\varphi^{0}z^{2n+1}=z\psi\,,$
where $\psi\in F[x,y,z]$ is a homogeneous polynomial of degree $2n$, and as in the previous case, we obtain that $\psi$ is also a solution. ∎
@@ 4. Proof of the main results
$\varphi_{x}$ $=z\cdot p_{s}(xz+y^{2},z)\,,$
$\varphi_{y}$ $=2y\cdot p_{s}(xz+y^{2},z)\,.$
in the linear PDE (5), one verifies that $\varphi=p(xz+y^{2},z)$ is a solution. Reciprocally, suppose that $\varphi\in F[x,y,z]$ is a solution of the equation (5). We shall prove that $\varphi$ has the form $\varphi=p(xz+y^{2},z)$, where $p$ is a polynomial in $F[t_{1},t_{2}]$. By Lemma 3, we can assume that $\varphi\in F[x,y,z]$ is a homogeneous polynomial. More precisely, we shall prove by recurrence the following:
@@
$\varphi=\left\{\begin{array}[]{ccl}a(xz+y^{2})+z\psi&\mbox{if}&\deg\varphi=2\,,\\ &&\\ z\psi_{1}&\mbox{if}&\deg\varphi=1\,,\end{array}\right.$
@@
Let us suppose that (18) holds for homogeneous polynomial solutions of degree $2n$ or $12n-$ 1. If $\deg\varphi=2n+2$, then by Proposition 4 ($a$), we obtain that $\varphi=a(xz+y^{2})^{2n+2}+z\psi$, where $\psi$ is a solution of the equation (5). By the inductive hypothesis, we have
@@
If $\deg\varphi=2n+1$, then by Proposition 4 ($b$), we obtain that $\varphi=z\psi$, where $\psi$ is a solution of the equation (5) with $\deg\psi=2n$. By the inductive hypothesis, one has
@@
which completes the proof. ∎
Proof of the Theorem 2.
One implication is straightforward. Suppose that $\det J(f,g,h)$ is a nonzero constant. By Proposition 1, $\varphi$ satisfies the linear PDE (5). By Theorem 1, the solution of the equation (5) has the form $\varphi=p(xz+y^{2},z)$ where $p$ is a polynomial in $F[t_{1},t_{2}]$. Finally, by Proposition 2, $(f,g,h)$ is an automorphism and its inverse is given in (2). ∎
@@ 5. Conclusions
Theorem 2 says that the automorphisms of the form (2) satisfy the well know Jacobian conjecture via the approach of solving first order partial differential equations. A further problem is whether those automorphisms might be naturally extended on polynomial rings with more than three variables satisfying the aforementioned result.

### v2

@@ 3. Statement of the main results
(11) $\varphi=\phi^{2n}+\phi^{2n-1}z+\cdots+\phi^{1}z^{2n-1}+\phi^{0}z^{2n}\,$,
where $\phi^{2n-k}\in F[x,y]$ is a homogeneous polynomial of degree 2nk2n-k for $0\leq k\leq 2n$. By Lemma 4, we get the following equalities:
(12) $\phi^{2n}$ $=ay^{2n}\,\,\,\,\textnormal{and}\,\,\,\,\phi^{1}=bx$
(13) $2y\phi^{2n-k-1}_{x}$ $=\phi^{2n-k}_{y}\,\,\,\textnormal{para todo}\,\,\,k=0,1,\dots,2n-2\,.$
We claim that there are constants $a,a^{2n-k}_{k-1-j}$, with $1\leq k\leq n$ y $0\leq j\leq k-1$, such that:
(14) $\phi^{2n-k}$ $=a\binom{n}{k}y^{2n-2k}x^{k}+\sum_{j=0}^{k-1}a^{2n-k}_{k-1-j}y^{2n-2k+1+j}x^{k-1-j}$
Indeed, taking $k=0$, from the equalities (12) and (13), we find that:
$2y\phi^{2n-1}_{x}=\phi^{2n}_{y}=2nay^{2n-1}\,$,
so that $\phi_{x}^{2n-1}=nay^{2n-2}$. Hence, integrating with respect to $x$, we obtain
$\phi^{2n-1}=nay^{2n-2}x+a^{2n-1}_{0}y^{2n-1}\,$,
@@
$\phi^{2n-1}=a\binom{n}{1}y^{2n-2}x+a^{2n-1}_{0}y^{2n-1}\,$,
thus (14) is true for $k=1$. Now, suppose that we have (14). Taking the derivative with respect to $y$, we have
$\phi^{2n-k}_{y}=2a\binom{n}{k}(n-k)y^{2n-2k-1}x^{k}+\sum_{j=0}^{k-1}\,(2n-2k+1+j)a^{2n-k}_{k-1-j}y^{2n-2k+j}x^{k-1-j}$
By the equality (13), we get
$\phi^{2n-k-1}_{x}=a\binom{n}{k}(n-k)y^{2n-2k-2}x^{k}+\sum_{j=0}^{k-1}\frac{(2n-2k+1+j)a^{2n-k}_{k-1-j}}{2}y^{2n-2k-1+j}x^{k-1-j}\,.$
Integrating with respect to $x$, there exists a constant $a^{2n-k-1}_{0}\in F$ such that
$\phi^{2n-k-1}=a\binom{n}{k+1}y^{2n-2k-2}x^{k+1}+\sum_{j=0}^{k-1}a^{2n-k-1}_{k-j}y^{2n-2k-1+j}x^{k-j}+a^{2n-k-1}_{0}y^{2n-k-1}\,$,
where $a^{2n-k-1}_{k-j}=\frac{(2n-2k+1+j)a^{2n-k}_{k-1-j}}{2(k-j)}$ for all $0\leq j\leq k-1$. Thus, we obtain
$\phi^{2n-k-1}=a\binom{n}{k+1}y^{2n-2k-2}x^{k+1}+\sum_{j=0}^{k}a^{2n-k-1}_{k-j}y^{2n-2k-1+j}x^{k-j}\,$,
as wished, and this completes the recurrence.
Now, replacing the expressions (12) and (14) in the equation (11), we have
$\varphi$ $=\phi^{2n}+\sum_{k=1}^{n}\phi^{2n-k}z^{k}+\sum_{k=n+1}^{2n}\phi^{2n-k}z^{k}$
$=ay^{2n}+\sum_{k=1}^{n}\left(a\binom{n}{k}y^{2n-2k}x^{k}+\sum_{j=0}^{k-1}a^{2n-k}_{k-1-j}y^{2n-2k+1+j}x^{k-1-j}\right)z^{k}+\sum_{k=n+1}^{2n}\phi^{2n-k}z^{k}$
$=a\sum_{k=0}^{n}\binom{n}{k}y^{2n-2k}x^{k}z^{k}+\sum_{k=1}^{n}\left(\sum_{j=0}^{k-1}a^{2n-k}_{k-1-j}y^{2n-2k+1+j}x^{k-1-j}\right)z^{k}+\sum_{k=n+1}^{2n}\phi^{2n-k}z^{k}\,.$
From the last equality, we can write
$\varphi=a(xz+y^{2})^{n}+z\psi,$
where $\psi\in F[x,y,z]$ is a homogeneous polynomial of degree equal to 2n12n-1. By the linearity of the solutions of the equation (7), we deduce that $z\psi$ is a solution, hence $\psi$ is also a solution.
($b$) Let us suppose that:
(15) $\varphi=\phi^{2n+1}+\phi^{2n}z+\cdots+\phi^{1}z^{2n}+\phi^{0}z^{2n+1}\,$,
where $\phi^{2n+1-k}$ is a homogeneous polynomial in $F[x,y]$ of degree equal to 2n+1k2n+1-k, for all $0\leq k\leq 2n+1$. By Lemma 4, we get the following equalities:
(16) $\phi^{2n+1}$ $=ay^{2n+1}$
(17) $2y\phi^{2n-k}_{x}$ $=\phi^{2n+1-k}_{y}\,\,\,\textnormal{para todo}\,\,\,k=0,1,\dots,2n-1\,.$
As in the even case, we claim that there are constants $a,b_{k-j}^{2n-k}$, with $0\leq k\leq n-1$ and $0\leq j\leq k$, such that:
(18) $\phi^{2n-k}=P^{2n+1}_{k}ay^{2n-2k-1}x^{k+1}+\sum_{j=0}^{k}b^{2n-k}_{k-j}y^{2n-2k+j}x^{k-j}\,$,
where $P^{2n+1}_{k}=\frac{(2n+1)\cdot(2n+1-2)\cdots(2n+1-2k)}{2^{k+1}\cdot(k+1)!}$ for all index $0\leq k\leq n-1$.
Replacing $k=0$ in the equation (16) and joining with (17), we have
$2y\phi^{2n}_{x}=\phi^{2n+1}_{y}=(2n+1)ay^{2n}\,$,
hence $\phi_{x}^{2n}=\cfrac{(2n+1)}{2}\,\,ay^{2n-1}$. Integrating with respect to $x$, we obtain
$\phi^{2n}=\cfrac{(2n+1)}{2}\,ay^{2n-1}x+b^{2n}_{0}y^{2n}$
thus (18) is true for $k=0$. Suppose that we have (18). Taking the derivative with respect to $y$, we have
$\phi^{2n-k}_{y}=P^{2n+1}_{k}(2n-2k-1)ay^{2n-2k-2}x^{k+1}+\sum_{j=0}^{k}(2n-2k+j)b^{2n-k}_{k-j}y^{2n-2k+j-1}x^{k-j}\,.$
@@
$\phi^{2n-k-1}_{x}=\frac{P^{2n+1}_{k}(2n-2k-1)}{2}\,ay^{2n-2k-3}x^{k+1}+\sum_{j=0}^{k}\frac{(2n-2k+j)b^{2n-k}_{k-j}}{2}y^{2n-2k+j-2}x^{k-j}\,.$
Integrating with respect to $x$, we find a constant $b^{2n-k-1}_{0}\in F$ such that $\phi^{2n-k-1}$ is equal to
$P^{2n+1}_{k+1}ay^{2n-2k-3}x^{k+2}+\sum_{j=0}^{k}\frac{(2n-2k+j)b^{2n-k}_{k-j}}{2(k-j+1)}y^{2n-2k+j-2}x^{k-j+1}+b^{2n-k-1}_{0}y^{2n-k-1}\,$,
taking into account that $P^{2n+1}_{k+1}=\frac{P^{2n+1}_{k}\cdot(2n-2k-1)}{2\cdot(k+2)}$. Defining
$b^{2n-k-1}_{k+1-j}=\frac{(2n-2k+j)b^{2n-k}_{k-j}}{2(k-j+1)}$
for all $0\leq j\leq k$, we get
$\phi^{2n-k-1}=P^{2n+1}_{k+1}\,ay^{2n-2k-3}x^{k+2}+\sum_{j=0}^{k+1}b^{2n-k-1}_{k+1-j}y^{2n-2k+j-2}x^{k-j+1}$
as wished, and this completes the recurrence. In particular, replacing k=n1k=n-1 in (18), we obtain:
$\phi^{n+1}=\frac{1\cdot 3\cdots(2n+1)}{2^{n}\cdot n!}ayx^{n}+\sum_{j=0}^{n-1}b^{n+1}_{n-1-j}y^{j+2}x^{n-1-j}\,.$
Considering the equation (17) for $k=n$, we get $2y\phi^{n}_{x}=\phi^{n+1}_{y}$, hence
(19) $2y\phi^{n}_{x}$ $=\frac{1\cdot 3\cdots(2n+1)}{2^{n-1}\cdot n!}ax^{n}+\sum_{j=0}^{n-1}(j+2)b^{n+1}_{n-1-j}y^{j+1}x^{n-1-j}\,.$
From this equality, we deduce that $a=0$, and consequently $\phi^{2n+1}=0$. Therefore,
$\varphi=z\phi^{2n}+z^{2}\phi^{2n-1}+\cdots+z^{2n}\phi^{1}+\phi^{0}z^{2n+1}=z\psi\,$,
where $\psi\in F[x,y,z]$ is a homogeneous polynomial of degree $2n$. Finally, arguing as in item ($a$), we deduce that $\psi$ is also a solution.
@@ 4. Proof of the main results
$\varphi_{x}$ $=z\cdot p_{t_{1}}(xz+y^{2},z)\,$,
$\varphi_{y}$ $=2y\cdot p_{t_{1}}(xz+y^{2},z)\,.$
in the linear PDE (7), one readily verifies that $\varphi=p(xz+y^{2},z)$ is a solution. Reciprocally, suppose that $\varphi\in F[x,y,z]$ is a solution of the equation (7). We shall prove that $\varphi$ has the form $\varphi=p(xz+y^{2},z)$, where $p$ is a polynomial in $F[t_{1},t_{2}]$. By Lemma 3, we can assume that $\varphi\in F[x,y,z]$ is a homogeneous polynomial. More precisely, we shall prove by recurrence the following:
@@
[…]
@@
Let us suppose that (20) holds for homogeneous polynomial solutions of degree $2n$ or 2n12n-1. If $\deg\varphi=2n+2$, then by Proposition 4 ($a$), we obtain that $\varphi=a(xz+y^{2})^{2n+2}+z\psi$, where $\psi$ is a solution of the equation (7). By the inductive hypothesis, we have
@@
Therefore, $\varphi$ has the form $\varphi=p(xz+y^{2},z)$, as wished. On the other hand, if $\deg\varphi=2n+1$, then by Proposition 4 ($b$), we obtain that $\varphi=z\psi$, where $\psi$ is a solution of the equation (7) with $\deg\psi=2n$. By the inductive hypothesis, one has
@@
thus $\varphi$ has the form $\varphi=p(xz+y^{2},z)$, and this completes the proof. ∎
Proof of the Corollary 3.
One implication is straightforward. Suppose that the determinant of $J(f,g,h)$ is a nonzero constant. By Proposition 1, $\varphi$ satisfies the linear PDE (7). By Theorem 1, the solution of the equation (7) has the form $\varphi=p(xz+y^{2},z)$, where $p$ is a polynomial in $F[t_{1},t_{2}]$. Finally, by Proposition 2, $(f,g,h)$ is an automorphism and its inverse is given in (3). ∎
@@ 5. Conclusions
In Corollary 3, we show that the Nagata automorphisms of the form (2) satisfy the well-known Jacobian conjecture. The proof is based on algebraic techniques for solving first-order linear partial differential equations. This result suggests a new strategy for approaching this longstanding open problem.